(2xy+3y2)dx-(2xy+X2)dy=0

1
Expert's answer
2021-12-23T06:26:44-0500
"- \frac{{dy}}{{dx}}( {x^2} + 2xy) +2xy+ 3{y^2} = 0"


Let "y(x) = xv(x)", which gives "\dfrac{dy}{dx}=v+x\dfrac{dv}{dx}" :

"-(v+x\dfrac{dv}{dx})( x^2 + 2x^2v) +2x^2v+3x^2v^2 = 0"

Simplify:


"x^2(-x\dfrac{dv}{dx}(1+2v)-v-2v^2+2v+3v^2)=0"

"\dfrac{1+2v}{v(v+1)}dv=\dfrac{dx}{x}"

"(\dfrac{1}{v}+\dfrac{1}{v+1})dv=\dfrac{dx}{x}"

Integrate


"\int(\dfrac{1}{v}+\dfrac{1}{v+1})dv=\int\dfrac{dx}{x}"

"\ln(|v|)+\ln(|v+1|)=\ln(|x|)+\ln c_1"

"v(v+1)=c_1x"

"xv(xv+x)=c_1x^3"

"y^2+xy=c_1x^3"